OpenStudy (hoopcoach99):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (anonymous):

m = -3/2 "slope" m = (y1-y0)/(x1-x0) "rise ofer run" so -3/2 = ((6-(-3))/(x-10) solve for x

OpenStudy (anonymous):

you must know the formulas by heart.. even better is to know why they work work... then it these become easy